GSIA, 45-733 Carnegie Mellon University Probability and Statistics I Spring, 2001-2002

Syllabus

Lectures  section M: TR 6pm-7:50, GSIA 240

section F: TR 8pm-9:50, GSIA 152

Instructor  Bill Vogt

HBH 2116D, 268-1843 wilibear@andrew.cmu.edu

Office Hours T 5:00-6:00

R 5:00-6:00

Grader  Puneet Goel, pgoel@andrew.cmu.edu
Textbook  Newbold, Paul (1995)

Statistics for Business & Economics, 4th ed Upper Saddle River, NJ: Prentice-Hall

Website  http://www.andrew.cmu.edu/course/45-733/index.htm

1 Course Objectives

Our objective will be to build understanding of elementary tools of statistical estimation and inference.

2 Prerequisites

This is the first course in a two course series in probability and statistics. The student is expected to be competent in algebra at the undergraduate level.

3 Grading

The grade will be determined by performance on homework assignments, a midterm (33%), and a final exam. Each will have equal weight. You are responsible for the content of the lectures, including any handouts, and chapters 1-9 in the text. The lectures cover material very similar to that presented in chapters 1-9 of the text.

4 Various Class Policies

1. Policy on Cooperation

• You may cooperate as much as you like on homework assignments; however, each individual must submit a separate assignment, and each individual will be separately graded on that assignment.

• You may not cooperate while taking exams.

2. Policy on Aids During Exams

• All exams are open book. You may use text, notes, calculators, computers, reference materials, etc.

• You may not communicate or cooperate with anyone on the exam.

• Please do not use the open book policy as a substitute for studying. If you need to look in your book, notes, etc for instructions on HOW TO DO the exam, you will surely run out of time.

3. Policy on Format of Assigned Work

• All work must be legible. Illegible is the equivalent of completely wrong.

• Spelling, grammar, style, etc do not “count” per se. However, anything which I (or the grader) cannot understand is wrong. Poor spelling, grammar, style, etc are often confusing.

4. Policy on Lateness

• Due dates for assignments are in the syllabus in the schedule section. Unless the instructor announces otherwise, these dates are binding.

• Late assignments are not accepted without either prior arrangement or compelling and verifiable reason.

• Assignments and homeworks will be returned to you in class. You are responsible for keeping them to compare against your final grade. I will use whatever grade I have recorded for your work (even if you think there is an error) unless you can show me the graded work so that I can see the error.

• You are responsible for collecting your graded work. Claims by students similar to “I turned in homework #2, but I don’t have a copy of it because you did not return it to me!” must be made within one week of the return date of the assignment. If you fail to make the complaint in time, I will use whatever grade I have in my records.

Syllabus 45-733 Spring 2001/2 5

5 Schedule

The class schedule below is tentative and likely will not be followed exactly.

Date Material, Book Sections Work
T Jan 15 Introduction, Chap 1,2
R Jan 17 Probability, Chap 3
T Jan 22 Discrete Random Variables, Chapter 4
R Jan 24 Discrete Random Variables, Chapter 4  HWK 1 due
T Jan 29 Continuous Random Variables, Chapter 5
R Jan 31 Continuous Random Variables, Chapter 5 HWK 2 due
T Feb 5 Sampling Distributions, Chapter 6  HWK 3 due
R Feb 7 In-class Midterm
T Feb 12 Point Estimation, Chapter 7
R Feb 14 Interval Estimation, Chapter 8  HWK 4 due
T Feb 19 Interval Estimation, Chapter 8
R Feb 21 Hypothesis Testing, Chapter 9  HWK 5 due
T Feb 26 Hypothesis Testing, Chapter 9
R Feb 28 In-class final